OpenStudy (sachin00007):

Find dy/dx of y=3tan^4x

OpenStudy (satan ):

Apply chain rule.

OpenStudy (sachin00007):

could you pls help me to solve it :) @Satan 

OpenStudy (anonymous):

\[\begin{align*}\frac{dy}{dx}&=3\frac{d}{dx}\left[\tan^4x\right]\\ &=3\left(4\tan^3x\right)\frac{d}{dx}\left[\tan x\right]\\ &=3\left(4\tan^3x\right)\left(\sec^2x\right)\\ &=12\tan^3x\sec^2x\end{align*}\]